OpenStudy (anonymous):

1. The length of one leg of an isosceles right triangle is 3 ft. What is the perimeter of the triangle? 2. Angle A is equal to 33.69 degrees. Given that the horizontal distance from the hoop to the free-throw line (at point A) is 15 feet, how tall is the hoop? Round the answer to the nearest whole foot. 3. A 7-foot ladder is leaned against a building in such a way that the bottom of the ladder is 4 feet from the base of the wall. Find the angle formed between the ladder and the ground. i will give a medal thingy to whoever helps :)

OpenStudy (sofialearner101):

omg i have no idea

OpenStudy (sofialearner101):

sorry:(

OpenStudy (sofialearner101):

all i found was The other leg must be 3. The hypotenuse is 3sqrt(2) Perimeter = 3+3+3sqrt(2) = 6+3sqrt(2) ft

OpenStudy (sofialearner101):

but idk if that helps u